Circular No. 8803

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1280 SCORPII = NOVA SCORPII 2007 Reports of apparently independent discoveries of a possible nova have

been received by the Central Bureau: H. Yamaoka, Kyushu University, communicates the discovery of a new object (mag 9.9) by Yuji Nakamura (Kameyama, Mie, Japan) on two unfiltered CCD images taken on Feb. 4.8624 UT with a 135-mm camera lens, with the position of the variable measured to be α = 16^{h}57^{m}41^{s}.0, δ = −32^{o}20′34′′ (equinox 2000.0, uncertainty ∼ 30′′); nothing was detected at this position with the same equipment on Jan. 29.8669 (limiting mag 12.0) or Feb. 2.8662 (limiting mag 11.0). S. Nakano, Sumoto, Japan, reports the discovery by Yukio Sakurai (Mito, Ibaraki-ken) of the same star at mag 9.4 on two 10-s CCD exposures taken on Feb. 4.854 (correction to time given on CBET 834) using a Fuji FinePix S2 Digital Camera (+ Nikon 180-mm f/2.8 lens), the position measured to be α = 16^{h}57^{m}40^{s}.91, δ = −32^{o}20′36′′.4; nothing is visible on a CCD frame taken by Sakurai on Feb. 2.8. Nakano adds that he has measured the variable’s position from one of Sakurai’s images, yielding position end figures 41^{s}.24, 36′′.5. Nakano also writes that K. Kadota has confirmed the variable at mag 8.9 on an unfiltered CCD image taken on Feb. 5.818 with a 0.25-cm f/5 reflector, providing position end figures 41^{s}.20, 35′′.8.

H. Naito and S. Narusawa, Nishi-Harima Astronomical Observatory (NHAO), report that they obtained a low-dispersion spectrum (range 410-670 nm; resolution 1600 at Hα) of N Sco 2007 on Feb. 5.87 UT with the 2.0-m NAYUTA telescope (+ MALLS). The spectrum shows a smooth continuum with many lines (Hα, Hβ, He I 492- and 501.5-nm, and Fe II 516.3nm) with P-Cyg profiles, which suggests that the variable is indeed a classical nova caught near maximum light. The FWHM of the Hα emission is 400 km/s, and its absorption minimum is blueshifted by 480 km/s from the emission peak (measured by a Gaussian fitting).

N. N. Samus, Institute of Astronomy, Russian Academy of Sciences, advises the Bureau that the nova reported above has been given the designation V1280 Sco.

COMETS 184P/LOVAS AND 185P/PETRIEW Comet P/1986 W1 = 2007 A1 (cf. IAUC 8791) has been assigned the

permanent number 184P, and comet P/2001 Q2 = 2007 A3 (cf. IAUC 8795) has been assigned the number 185P (cf. MPC 58732).

2007 February 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green